OpenStudy (anonymous):

solve equation logx – 116 = 4

OpenStudy (anonymous):

Are u sure the equation is correct ?

OpenStudy (anonymous):

after the log the x-1 is small and then it's 16-4

OpenStudy (anonymous):

\[\log_x-16=4\]?

OpenStudy (anonymous):

Oh ,u mean Log(x-116)=4 ,or log(x) - 116=4

OpenStudy (anonymous):

I guess it's like log(x-1)16-4

OpenStudy (anonymous):

\[\log_{x-1}-16=4\]?

OpenStudy (anonymous):

oh yeah I 'm sorry after the 16 it is an equal sign

OpenStudy (anonymous):

if that,the answer is3 or-1

OpenStudy (anonymous):

\[\log_{x-1}16=4\]\[4^{x-1}=16\]\[4^{x-1}=4^{2}\]\[x-1=2\]\[x=3\] Why -1?\[4^{-2}={1\over16}\]

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

oh,i was wrong...